# On the order of lazy cellular automata

### v2

@@ Abstract
We study the most elementary family of cellular automata defined over an arbitrary group universe $G$ and an alphabet $A$: the lazy cellular automata, which act as the identity on configurations in $A^{G}$, except when they read a unique active transition $p\in A^{S}$, in which case they write a fixed symbol $a\in A$. As expected, the dynamical behavior of lazy cellular automata is relatively simple, yet subtle questions arise since they completely depend on the choice of $p$ and $a$. In this paper, we investigate the order of a lazy cellular automaton $\tau:A^{G}\to A^{G}$, defined as the cardinality of the set $\{\tau^{k}:k\in\mathbb{N}\}$. In particular, we establish a general upper bound for the order of $\tau$ in terms of $p$ and $a$, and we prove that this bound is attained when $p$ is a quasi-constant pattern.
@@ 1 Introduction
As expected, the dynamical behavior of a lazy CA is relatively simple, yet subtle questions arise because their evolution depends entirely on the choice of $p$ and $a$. In a certain sense, lazy cellular automata are even more elementary than the well-known elementary cellular automata (ECA) studied by Wolfram [14], where complex behavior already emerges. Since there are only 256 ECA, a complete case-by-case computational analysis is feasible, whereas there are infinitely many lazy CAs (over an infinite group universe), as their neighborhood size can be arbitrarily large. Despite this, we believe that a deep understanding of lazy CAs is possible and may lead to new insights into broader families of cellular automata.
@@
Besides being an important algebraic concept, the order also captures part of the dynamical behaviour of a cellular automaton. For example, it was shown by Kůrka [12, Theorem 4] that a one-dimensional cellular automaton (i.e, when the underlying universe is the group of integers $\mathbb{Z}$) is of finite order if and only if it is equicontinuous.
@@
In this paper, we establish several results on the order of lazy CAs in the general setting of an arbitrary group universe $G$. In Section 2, we show that for any lazy cellular automaton $\tau:A^{G}\to A^{G}$, if $\textup{ord}\left(\tau\right)$ is finite, then $\tau$ has period $1$, which means that there is $n\in\mathbb{N}$ such that $\tau^{n}=\tau^{n+1}$. Moreover, in Theorem 2, we derive a general upper bound for $\textup{ord}\left(\tau\right)$ in terms of its unique active transition $p\in A^{S}$. As a consequence, in Corollary 3, we provide a sufficient condition on $p\in A^{S}$ that guarantees that $\tau$ is idempotent.
In Section 3, we determine the order of a lazy CA whose unique active transition is a quasi-constant pattern, which is a substantial generalization of Theorem 2 in [3]. In order to state our result, we introduce the following notation about words on groups. A word on $S\subseteq G$ of length $n$ is simply an element of the Cartesian power $S^{n}:=\{(s_{1},\dots,s_{n}):s_{i}\in S\}$. We consider the evaluation function $\theta$ from words on $S\subseteq G$ to elements of $G$ given by $\theta(s_{1},\dots,s_{n}):=s_{1}\dots s_{n}$. We say that $v$ is a subword of $w=(s_{1},\dots,s_{n})$, denoted by $v\sqsubseteq w$, if $v=(s_{i},s_{i+1},\dots,s_{j})$ for some $i\leq j$.
Theorem 1.
Let $\tau:A^{G}\to A^{G}$ be a lazy cellular automaton with unique active transition $p\in A^{S}$ and writing symbol $a\in A^{S}\setminus\{p(e)\}$. Assume that $p$ is quasi-constant with non-constant element $r\in S$.
  1. 1.
    If $a\neq p(s)$ for all $s\in S$, then $\textup{ord}\left(\tau\right)=2$.
  2. 2.
    $\textup{ord}\left(\tau\right)=\min\{n\geq 2:r^{n}\in S\}.$
  3. 3.
    If $r=e$ and $a=p(s)$ for all $s\in S\setminus\{e\}$, then $\textup{ord}\left(\tau\right)$ is finite if and only if there exists $n\geq 2$ such that for all words $w\in(S\setminus\{e\})^{n-1}$ there exists a subword $v\sqsubseteq w$ such that $\theta(v)^{-1}\in S$. In such case, the order of $\tau$ is the minimum $n$ satisfying this property.
As an easy consequence of Theorem 1 we find that, in contrast with the dichotomy obtained in [4] when the neighborhood $S\subseteq\mathbb{Z}$ is an interval, for every $n\geq 2$, there exists a lazy cellular automaton $\tau:A^{G}\to A^{G}$ such that $\textup{ord}\left(\tau\right)=n$.
@@ 2 The order of lazy CA
We assume that the alphabet $A$ has at least two different elements 0 and $1$. A pattern is a function $p\in A^{S}$, where $S$ is a finite subset of the group universe $G$.
@@
In such case, we say that $p$ is the unique active transition of $\tau$ and $\mu(p)\in A\setminus\{p(e)\}$ is the writing symbol of $\tau$.
@@
Let $\mu:A^{S}\to A$ be the minimal local map of $\tau$. By Definition 1, there is $g\in G$ such that $(g\cdot x)|_{S}=p$ if and only if
$\tau(x)(g)=\mu((g\cdot x)|_{S})\neq p(e)=(g\cdot x)(e)=x(g).$
The result follows. ∎
@@
Let $\tau:A^{G}\to A^{G}$ be a lazy CA with unique active transition $p\in A^{S}$ and writing symbol $a\in A$. If there exists $x\in A^{G}$ such that $\tau(x)|_{S}=p$, then $(s\cdot x)|_{S}=p$ for some $s\in S\setminus\{e\}$ such that $p(s)=a\neq p(e)$.
@@
$p(s)=\tau(x)(s)=\mu((s\cdot x)|_{S}),\quad\forall s\in S,$
@@
If $x|_{S}=p$, then $p(e)=\mu(x|_{S})=a$ is a contradiction. Hence, $x|_{S}\neq p$ and $p(e)=\mu(x|_{S})=x(e)$. Suppose that $(s\cdot x)|_{S}\neq p$ for all $s\in S\setminus\{e\}$. Then,
$p(s)=\mu((s\cdot x)|_{S})=(s\cdot x)(e)=x(s),$
which contradicts that $x|_{S}\neq p$. Therefore, there must exist $s\in S\setminus\{e\}$ such that $(s\cdot x)|_{S}=p$. This implies that $p(s)=\mu((s\cdot x)|_{S})=a\neq p(e)$. ∎
@@
If $\textup{ord}\left(\tau\right)<\infty$, let $m\in\mathbb{N}$ and $n\in\mathbb{Z}_{+}$ be as small as possible such that $\tau^{m}=\tau^{m+n}$. Then, $m$ and $n$ are referred to as the index and period of $\tau$, respectively. Note that $\textup{ord}\left(\tau\right)=m+n$.
@@
Let $\mu:A^{S}\to A$ be the minimal local map of $\tau$ with unique active transition $p\in A^{S}$ and writing symbol $a:=\mu(p)\in A$. If $\textup{ord}\left(\tau\right)<\infty$, let $m$ and $n$ be the index and period of $\tau$, so $\tau^{m}=\tau^{m+n}$. Lemma 3 (1.) implies that for all $x\in A^{G}$
$\textup{supp}_{a}\left(\tau^{m}(x)\right)=\textup{supp}_{a}\left(\tau^{m+i}(x)\right),\quad\forall i\in\{1,\dots,n\}.$
@@
$\textup{ord}\left(\tau\right)>n\iff\tau^{j-1}\neq\tau^{j},\quad\forall j\in{1,\ldots,n}.$
Furthermore,
$\textup{ord}\left(\tau\right)=\min\{n\geq 2:\tau^{n-1}=\tau^{n}\}.$
We define the product of two subsets $S,K\subseteq G$ as $SK:=\{sk:s\in S,k\in K\}$ and the inverse as $S^{-1}:=\{s^{-1}:s\in S\}$. In the following result we provide a general upper bound for the order of a lazy cellular automaton in terms of its unique active transition.
@@
Let $\tau:A^{G}\to A^{G}$ be a lazy CA with minimal neighborhood $S\subseteq G$, unique active transition $p\in A^{S}$ and writing symbol $a\in A\setminus\{p(e)\}$. For any $b\in A$, we define
$S_{b}:=p^{-1}\{b\}=\{s\in S:p(s)=b\}.$
Then, $\textup{ord}\left(\tau\right)$ is at most the minimum $n\geq 2$ such that for every word $(s_{1},\ldots,s_{n-1})\in(S_{a})^{n-1}$ there exist $1\leq i\leq j\leq n-1$ satisfying at least one of the following:
@@
    $(s_{j}\cdots s_{i})^{-1}\in S_{b}^{-1}S_{a}$, for some $b\in A\setminus\{a\}$;
@@
    $(s_{j}\cdots s_{i})^{-1}\in S_{b_{1}}^{-1}S_{b_{2}}$, for some $b_{1},b_{2}\in A\setminus\{a\}$, with $b_{1}\neq b_{2}$.
@@
We will show that for any $n\geq 2$, if $\tau^{n-1}\neq\tau^{n}$, then there exists a word $(s_{1},\ldots,s_{n-1})\in(S_{a})^{n-1}$ such that (1.) and (2.) do not hold for all $1\leq i\leq j\leq n-1$, so the result follows from Corollary 2.
If $\tau^{n-1}\neq\tau^{n}$, then, by Corollary 1, there exists $x\in A^{G}$ such that $p$ appears in $\tau^{n-1}(x)$. By the $G$-equivariance of cellular automata (i.e., $\tau(g\cdot x)=g\cdot\tau(x)$, for all $g\in G$, see [6, Prop. 1.4.4]), we may assume that $\tau^{n-1}(x)|_{S}=p$. Applying iteratively Lemmas 1 and 2 yields the existence of a word $(s_{1},\ldots,s_{n-1})\in(S_{a})^{n-1}$ such that
$\big((s_{j}\cdots s_{1})\cdot\tau^{(n-1)-j}(x)\big)|_{S}=p,\quad\forall j\in\{1,\ldots,n-1\}.$
It follows from the definition of $S_{b}$ that for all $g\in G$, $y\in A^{G}$,
$(g\cdot y)|_{S}=p\quad\iff\quad S_{b}g\subseteq\textup{supp}_{b}\left(y\right),\quad\forall b\in A,$
Hence, for all $b\in A$,
$S_{b}(s_{j}\cdots s_{1})$ $\subseteq\textup{supp}_{b}\left(\tau^{(n-1)-j}(x)\right),\quad\forall j\in\{1,\ldots,n-1\},$
Now, Lemmas 3 (1.) and (2.) imply that for all $k\in\{1,\dots,n-1\}$,
@@
Finally, (1) is equivalent to $(s_{j}\cdots s_{i})^{-1}\notin S_{b}^{-1}S_{a}$ for all $1\leq i\leq j\leq n-1$ and $b\in A\setminus\{a\}$, while (2) is equivalent to $(s_{j}\cdots s_{i})^{-1}\notin S_{b_{1}}^{-1}S_{b_{2}}$, for all $1\leq i\leq j\leq n-1$, and all $b_{1},b_{2}\in A\setminus\{a\}$, $b_{1}\neq b_{2}$. The result follows.
@@
A counterexample of the converse of Corollary 3 was found by María G. Magaña-Chávez in private communication. Consider the lazy CA $\tau:A^{\mathbb{Z}}\to A^{\mathbb{Z}}$ with minimal neighborhood $S=\{-3,-1,0,3,4\}$, unique active transition $p=11010\in A^{S}$ and writing symbol $a=1\in A$. Then $S_{0}=\{0,4\}$ and $S_{1}=\{-3,-1,3\}$, so none of the conditions of Corollary 3 holds. However, direct computations show that $\tau$ is idempotent.

### v3

@@ Abstract
We study the most elementary family of cellular automata defined over an arbitrary group universe $G$ and an alphabet $A$: the lazy cellular automata, which act as the identity on configurations in $A^{G}$, except when they read a unique active transition $p\in A^{S}$, in which case they write a fixed symbol $a\in A$. As expected, the dynamical behavior of lazy cellular automata is relatively simple, yet subtle questions arise since they completely depend on the choice of $p$ and $a$. In this paper, we investigate the order of a lazy cellular automaton $\tau:A^{G}\to A^{G}$, defined as the cardinality of the set $\{\tau^{k}:k\in\mathbb{N}\}$. In particular, we establish a general upper bound for the order of $\tau$ in terms of the fibers of the pattern $p$, and we prove that this bound is attained when $p$ is a quasi-constant pattern.
@@ 1 Introduction
As expected, the dynamical behavior of a lazy CA is relatively simple, yet subtle questions arise because their evolution depends entirely on the choice of $p$ and $a$. In a certain sense, lazy cellular automata are even more elementary than the well-known elementary cellular automata (ECA) studied by Wolfram [14], where complex behavior already emerges. Since there are only 256 ECA, a complete case-by-case computational analysis is feasible, whereas there are infinitely many lazy CAs (over an infinite group universe), as their neighborhood size can be arbitrarily large. Despite this, a deep understanding of lazy CAs is possible and may lead to new insights into broader families of cellular automata.
@@
Besides being an important algebraic concept, the order also captures part of the dynamical behavior of a cellular automaton. For example, it was shown by Kůrka [12, Theorem 4] that a one-dimensional cellular automaton (i.e, when the underlying universe is the group of integers $\mathbb{Z}$) is of finite order if and only if it is equicontinuous.
@@
In this paper, we establish several results on the order of lazy CAs in the general setting of an arbitrary group universe $G$. In Section 2, we show that for any lazy cellular automaton $\tau:A^{G}\to A^{G}$, if $\textup{ord}\left(\tau\right)$ is finite, then $\tau$ has period $1$, which means that there is $n\in\mathbb{N}$ such that $\tau^{n}=\tau^{n+1}$. Moreover, in Theorem 1, we derive a general upper bound for $\textup{ord}\left(\tau\right)$ which depends purely on combinatorial properties of the fibers $S_{a}:=p^{-1}\{a\}\subseteq S$ of of its unique active transition $p\in A^{S}$. As a consequence, in Corollary 3, we provide a sufficient condition on $p\in A^{S}$ that guarantees that $\tau$ is idempotent.
In Section 3, we determine the order of a lazy CA whose unique active transition is a quasi-constant pattern, which is a substantial generalization of Theorem 2 in [3].
As an easy consequence of Theorem 2 we find that, in contrast with the dichotomy obtained in [4] when the neighborhood $S\subseteq\mathbb{Z}$ is an interval, for every $n\geq 2$, there exists a lazy cellular automaton $\tau:A^{\mathbb{Z}}\to A^{\mathbb{Z}}$ such that $\textup{ord}\left(\tau\right)=n$.
@@ 2 The order of lazy CA
We assume that the alphabet $A$ has at least two different elements 0 and $1$. A pattern is a function $p\in A^{S}$, where $S$ is a finite subset of the group universe $G$. We denote patterns by $p=p(s_{1})p(s_{2})\dots p(s_{n})$, where we fix an order on the set $S=\{s_{1},s_{2},\dots,s_{n}\}$.
@@
In such case, we say that $p$ is the unique active transition of $\tau$ and $a:=\mu(p)\in A\setminus\{p(e)\}$ is the writing symbol of $\tau$.
@@
By Definition 1, there is $g\in G$ such that $(g\cdot x)|_{S}=p$ if and only if
$\tau(x)(g)=\mu((g\cdot x)|_{S})\neq p(e)=(g\cdot x)(e)=x(g),$
where $\mu:A^{S}\to A$ is the minimal local map of $\tau$. The result follows. ∎
@@
Let $\tau:A^{G}\to A^{G}$ be a lazy CA with unique active transition $p\in A^{S}$ and writing symbol $a\in A$. If there exists $x\in A^{G}$ such that $\tau(x)|_{S}=p$, then $(s\cdot x)|_{S}=p$ for some $s\in S_{a}$.
@@
$p(s)=\tau(x)(s)=\mu((s\cdot x)|_{S}),\quad\forall s\in S,$ (1)
@@
If $x|_{S}=p$, then $p(e)=\mu(x|_{S})=a$ is a contradiction with Definition 1. Hence, $x|_{S}\neq p$ and $p(e)=\mu(x|_{S})=x(e)$ by the hypothesis (1) with $s=e$. Suppose that $(s\cdot x)|_{S}\neq p$ for all $s\in S\setminus\{e\}$. Then, by hypothesis (1),
$p(s)=\mu((s\cdot x)|_{S})=(s\cdot x)(e)=x(s),\quad\forall s\in S\setminus\{e\},$
which contradicts that $x|_{S}\neq p$. Therefore, there must exist $s\in S\setminus\{e\}$ such that $(s\cdot x)|_{S}=p$. Again hypothesis (1) gives us that that $p(s)=\mu((s\cdot x)|_{S})=a$, so $s\in S_{a}$. ∎
@@
If $\textup{ord}\left(\tau\right)<\infty$, let $m\in\mathbb{N}$ and $n\in\mathbb{Z}_{+}$ be as small as possible such that $\tau^{m}=\tau^{m+n}$. Then, $m$ and $n$ are referred to as the index and period of $\tau$, respectively. Note that $\textup{ord}\left(\tau\right)=m+n$, as depicted in the following diagram:
$\mathrm{id}$$\tau$$\cdots$$\tau^{m}$$\cdots$$\tau^{m+n-1}$}
@@
Let $\mu:A^{S}\to A$ be the minimal local map of $\tau$ with unique active transition $p\in A^{S}$ and writing symbol $a:=\mu(p)\in A$. If $\textup{ord}\left(\tau\right)<\infty$, let $m$ and $n$ be the index and period of $\tau$, so $\tau^{m}=\tau^{m+n}$. Lemma 3 (1.) implies that for all $x\in A^{G}$,
$\textup{supp}_{a}\left(\tau^{m}(x)\right)\subseteq\textup{supp}_{a}\left(\tau^{m+1}(x)\right)\subseteq\dots\subseteq\textup{supp}_{a}\left(\tau^{m+n}(x)\right).$
However, since $\tau^{m}=\tau^{m+n}$, then $\textup{supp}_{a}\left(\tau^{m}(x)\right)=\textup{supp}_{a}\left(\tau^{m+n}(x)\right)$, so
$\textup{supp}_{a}\left(\tau^{m}(x)\right)=\textup{supp}_{a}\left(\tau^{m+1}(x)\right),\quad\forall x\in A^{G}.$
@@
$\textup{ord}\left(\tau\right)=\min\{n\geq 2:\tau^{n-1}=\tau^{n}\},$
where the minimum exists if and only if $\textup{ord}\left(\tau\right)<\infty$.
We define the product of two subsets $S,K\subseteq G$ as $SK:=\{sk:s\in S,k\in K\}$ and the inverse as $S^{-1}:=\{s^{-1}:s\in S\}$.
Lemma 4.
For all $g\in G$, $x\in A^{G}$, $p\in A^{S}$, it holds that
$(g\cdot x)|_{S}=p\quad\iff\quad S_{b}g\subseteq\textup{supp}_{b}\left(x\right),\ \forall b\in A.$
Proof.
Suppose that $(g\cdot x)|_{S}=p$, which is equivalent to $(x)(sg)=p(s)$, $\forall s\in S$. For any $b\in A$, if $s\in S_{b}$, then $x(sg)=p(s)=b$. This shows that $S_{b}g\subseteq\textup{supp}_{b}\left(x\right)$ for all $b\in A$.
Conversely, suppose that $S_{b}g\subseteq\textup{supp}_{b}\left(x\right)$, for all $b\in A$. Then, for all $s\in S$,
$sg\in S_{p(s)}g\subseteq\textup{supp}_{p(s)}\left(x\right).$
This means that $x(sg)=p(s)$, for all $s\in S$, and the result follows. ∎
In order to state the main theorem of this section, we introduce the following notation about words on groups. A word on $S\subseteq G$ of length $n$ is simply an element of the Cartesian power $S^{n}:=\{(s_{1},\dots,s_{n}):s_{i}\in S\}$, for $n\geq 1$ (we do not consider empty words). We consider the evaluation function $\theta$ from words on $S\subseteq G$ to elements of $G$ given by $\theta(s_{1},\dots,s_{n}):=s_{1}\dots s_{n}$. We say that $v$ is a subword of $w=(s_{1},\dots,s_{n})\in S^{n}$, denoted by $v\sqsubseteq w$, if $v=(s_{i},s_{i+1},\dots,s_{j})$ for some $i\leq j$. For integers $i,j\in\mathbb{Z}$ such that $i\leq j$, we consider the integer interval $i,j]:=\{k\in\mathbb{Z}:i\leq k\leq j\}$.
In the following result we provide a general upper bound for the order of a lazy cellular automaton in terms of a combinatorial condition satisfied by the fibers of its unique active transition.
@@
Let $\tau:A^{G}\to A^{G}$ be a lazy CA with minimal neighborhood $S\subseteq G$, unique active transition $p\in A^{S}$ and writing symbol $a\in A\setminus\{p(e)\}$. Then, $\textup{ord}\left(\tau\right)$ is at most the minimum $n\geq 2$ such that every word $w\in(S_{a}^{-1})^{n-1}$ has a subword $v\sqsubseteq w$ such that:
@@
    $\theta(v)\in S_{b}^{-1}S_{a}$, for some $b\in A\setminus\{a\}$; or
@@
    $\theta(v)\in S_{b_{1}}^{-1}S_{b_{2}}$, for some $b_{1},b_{2}\in A\setminus\{a\}$, with $b_{1}\neq b_{2}$.
@@
We will show that for any $n\geq 2$, if $\tau^{n-1}\neq\tau^{n}$, then there exists a word $w=(s^{-1}_{1},\ldots,s^{-1}_{n-1})\in(S^{-1}_{a})^{n-1}$ such that (1.) and (2.) do not hold for all subwords $v\sqsubseteq w$, so the result follows from Corollary 2.
If $\tau^{n-1}\neq\tau^{n}$, then, by Corollary 1, there exists $x\in A^{G}$ such that $p$ appears in $\tau^{n-1}(x)$. By the $G$-equivariance of cellular automata (i.e., $\tau(g\cdot x)=g\cdot\tau(x)$, for all $g\in G$, see [6, Prop. 1.4.4]), we may assume that $\tau^{n-1}(x)|_{S}=p$. Applying iteratively Lemmas 1 and 2 yields the existence of $s_{1},\dots,s_{n-1}\in S_{a}$ such that
$\big((s_{j}\cdots s_{1})\cdot\tau^{(n-1)-j}(x)\big)|_{S}=p,\quad\forall j\in[1,n-1].$
By Lemma 4 applied to the previous equality we obtain that for all $b\in A$,
$S_{b}(s_{j}\cdots s_{1})$ $\subseteq\textup{supp}_{b}\left(\tau^{(n-1)-j}(x)\right),\quad\forall j\in[1,n-1],$
Now, Lemmas 3 (1.) and (2.) imply that for all $k\in[1,n-1]$,
@@
Observe that condition (2) is equivalent to $(s_{j}\cdots s_{i})^{-1}=s_{i}^{-1}\dots s_{j}^{-1}\notin S_{b}^{-1}S_{a}$ for all $1\leq i\leq j\leq n-1$ and $b\in A\setminus\{a\}$, while condition (3) is equivalent to $(s_{j}\cdots s_{i})^{-1}=s_{i}^{-1}\dots s_{j}^{-1}\notin S_{b_{1}}^{-1}S_{b_{2}}$, for all $1\leq i\leq j\leq n-1$, and all $b_{1},b_{2}\in A\setminus\{a\}$, $b_{1}\neq b_{2}$. This means that no subword of $w:=(s_{1}^{-1},\dots,s_{n-1}^{-1})\in(S_{a}^{-1})^{n-1}$ is neither in $S_{b}^{-1}S_{a}$ nor in $S_{b_{1}}^{-1}S_{b_{2}}$, and the result follows. ∎
Example 3.
Consider the the lazy CA $\tau:A^{\mathbb{Z}}\to A^{\mathbb{Z}}$ with minimal neighborhood $S=\{-1,0,1\}$, unique active transition $p=101\in A^{S}$ and writing symbol $a=1\in A$ (this is ECA rule 236 used in Example 1). As the group $\mathbb{Z}$ uses additive notation, then condition (1.) of Theorem 1 involves the set
$S_{0}+S_{1}=\{0\}+\{-1,1\}=\{-1,1\}.$
Since the evaluation of any word of length $1$ over $S_{1}=\{1,-1\}$ is in $S_{0}+S_{1}$, Theorem 1 shows that the order of $\tau$ is at most order $2$. Since $\tau$ may not have order $1$ because it is not the identity function, then $\textup{ord}\left(\tau\right)=2$.
Example 4.
Consider the the lazy CA $\tau:A^{\mathbb{Z}}\to A^{\mathbb{Z}}$ with minimal neighborhood $S=\{-1,0,1\}$, unique active transition $p=001\in A^{S}$ and writing symbol $a=1\in A$ (this is ECA rule 206). Condition (1.) of Theorem 1 involves the set
$S_{0}+S_{1}=\{0,1\}+\{1\}=\{1,2\}.$
Since words of any length over $S_{1}=\{-1\}$ only evaluate to negative integers, then Theorem 1 does not give us an upper bound for the order of $\tau$. It may be easily seen that $\tau$ has indeed infinite order, as the configuration $x=\dots 000100\dots\in A^{\mathbb{Z}}$ never stabilizes.
Example 5.
Consider the the lazy CA $\tau:A^{\mathbb{Z}}\to A^{\mathbb{Z}}$ with minimal neighborhood $S=\{0,1,3\}$, unique active transition $p=01\_0\in A^{S}$ and writing symbol $a=1\in A$ (we use the symbol _\_ in the notation of $p$ to emphasize that $S$ is not an interval of integers). Condition (1.) of Theorem 1 involves the set
$S_{0}+S_{1}=\{0,-3\}+\{1\}=\{1,-2\}.$
The evaluation of the word of length $1$ over $S_{1}=\{-1\}$ is not in $S_{0}+S_{1}$, but the evaluation of the word of length $2$ over $S_{1}=\{-1\}$ is in $S_{0}+S_{1}$; hence, Theorem 1 shows that the order of $\tau$ is at most $3$. In Section 3 we will show that quasi-constant patterns such as p=01_0p=01\_0 always define lazy CA whose order achieve the upper bound given by Theorem 1.
@@
A counterexample of the converse of Corollary 3 was found by María G. Magaña-Chávez in private communication. Consider the lazy CA $\tau:A^{\mathbb{Z}}\to A^{\mathbb{Z}}$ with minimal neighborhood $S=\{-3,-1,0,3,4\}$, unique active transition $p=11010\in A^{S}$ and writing symbol $a=1\in A$. Then $S_{0}=\{0,4\}$ and $S_{1}=\{-3,-1,3\}$, so none of the conditions of Corollary 3 holds. However, $\tau$ is idempotent, as it may be checked that no pattern of $A^{S+S}$ evaluates to $p$ under the local rule of $\tau$ (see [3, Remark 3.3]).